OpenStudy (anonymous):

What is the 9th term of the geometric sequence where a1 = -5 and a6 = -5,120?

OpenStudy (anonymous):

You can first find the common ratio r by seeing the number of spaces apart the 2 sums are

OpenStudy (anonymous):

\[a_{n2}-a_{n1} = 6-1 = 5 \]

OpenStudy (anonymous):

So now we know \[a_6 = a_1r^5\]

OpenStudy (anonymous):

Solve for r and then use the first term a1 to solve for the nth term (9th) using your geometric formula

OpenStudy (dls):

Let the GP be a,ar,ar^2........ar^8 given that a1=a=-5 a6=(-5)r^5=-5120 r^5=1024=2^10 r=4 Therefore required term i.e a9 = a r^8 = > (-5) (4)^8

OpenStudy (anonymous):

\[a_6 = a_1r^5 \] \[-5,120 = (-5)r^5\] Dividing both sides by (-5)\[1024 = r^5 \] Taking the 5th root of both sides\[\sqrt[5]{1024} = r = 4\]

OpenStudy (anonymous):

\[a_{nth}= a_1(r^{n-1})\] \[a_9= (-5)(4^{9-1})\]